OpenStudy (anonymous):

the second angle of a triangle is three times the first and the third is 12 degrees less than twice the first. find the measures of the angles. I need help setting it up!

hero (hero):

m angle 1 = x m angle 2 = 3x m angle 3 = 2x - 12 Since it is a triangle, the three angles will add to get 180 degrees, therefore angle 1 + angle 2 + angle 3 = 180 degrees Thus: x + 3x + (2x - 12) = 180 x + 3x + 2x - 12 = 180 (1 + 3 + 2)x - 12 = 180 6x - 12 = 180 6x = 180 + 12 6x = 192 x = 192/6 x = 32 Now, insert the value of x back into the corresponding expressions given for each angle to find the measures of each.

OpenStudy (anonymous):

so angle 1=32 , angle 2=96, and angle 3=52 ?